### v1

# DiEC: Diffusion Embedded Clustering

Abstract
Deep clustering hinges on learning representations that are inherently clusterable. However, using a single encoder to produce a fixed embedding ignores the representation trajectory formed by a pretrained diffusion model across network hierarchies and noise timesteps, where clusterability varies substantially. We propose DiEC (Diffusion Embedded Clustering), which performs unsupervised clustering by directly reading internal activations from a pretrained diffusion U-Net.
DiEC formulates representation selection as a two-dimensional search over layer $\times$ timestep, and exploits a weak-coupling property to decompose it into two stages. Specifically, we first fix the U-Net bottleneck layer as the Clustering Middle Layer (CML), and then use Optimal Timestep Search (OTS) to identify the clustering-optimal timestep ($t^{*}$). During training, we extract bottleneck features at the fixed $t^{*}$ and obtain clustering representations via a lightweight residual mapping. We optimize a DEC-style KL self-training objective, augmented with adaptive graph regularization and entropy regularization to strengthen cluster structures. In parallel, we introduce a denoising-consistency branch at random timesteps to stabilize the representations and preserve generative consistency. Experiments show that DiEC achieves competitive clustering performance on multiple standard benchmarks.
1 Introduction
Figure 1: Overview of DiEC (Diffusion Embedded Clustering). Given an input image $x_{0}$, we inject noise using both a random timestep $t_{\mathrm{rand}}$ and a fixed clustering-optimal timestep (COT). The random-timestep branch preserves generative consistency via standard denoising, while the fixed-timestep branch extracts the Clustering Middle Layer(CML) for clustering-aware optimization. We apply residual decoupling on CML and jointly optimize a graph regularizer and a KL-based clustering objective to obtain more stable and discriminative clustering representations.
Clustering is a fundamental task in unsupervised learning, aiming to uncover latent class structure from unlabeled data, and has been widely used in applications such as image understanding vangansbeke2020scan, anomaly detection pang2021deepanomalyreview, and large-scale retrieval johnson2019billion. However, modern data are often high-dimensional, nonlinear, and corrupted by noise, making clustering increasingly dependent on high-quality and robust feature representations zhou2024csurdeepclustering. Traditional approaches typically rely on shallow features or simple distance metrics, and thus struggle to capture deep semantic structure in complex data bendavid2018clusteringwrong. Recent studies have therefore introduced deep models into clustering by jointly optimizing representation learning and clustering objectives, improving performance in challenging settings zhou2024csurdeepclustering. This line of research is commonly referred to as deep clustering and has been validated by representative methods such as Ren et al. ren2024tnnlsdeepclustering and Wei et al. wei2024overviewdeepclustering.
Recent progress in deep clustering under high-dimensional, nonlinear, and noisy settings has been largely driven by learning more semantic representations with deep networks, which in turn induces clearer cluster structures in the learned space zhou2024survey. Accordingly, much of the literature centers on the question of how to obtain clustering-friendly representations zhou2024survey. Existing approaches have broadly evolved along two complementary lines: (i) generative encoders, represented by autoencoders, variational autoencoders, and GANs, which learn low-dimensional embeddings via reconstruction or generation objectives xie2016dec; jiang2017vade; mukherjee2019clustergan; and (ii) discriminative self-supervised learning, represented by contrastive learning, which enforces consistency constraints to learn discriminative representations vangansbeke2020scan; caron2020swav. Despite the different training signals, both families typically share a common modeling assumption: the feature extractor naturally yields a single, most clustering-discriminative primary representation, on top of which cluster assignment and structure refinement are performed xie2016dec; zhou2024survey.
We observe that the clusterability of internal representations in diffusion denoising networks is not random; instead, it varies systematically with both the noise scale and the network hierarchy. Along the noise-scale dimension, clustering performance typically follows a unimodal trend, improving at moderate noise levels and degrading as the noise becomes too strong. Along the hierarchy dimension, representations around the bottleneck strike a more favorable balance between stability and separability. Moreover, across datasets, bottleneck activations are generally easier to access and more stable, whereas representations after the bottleneck often incur higher extraction cost with limited gains, resulting in a lower overall cost–benefit trade-off.
Motivated by the above patterns, we propose Diffusion Embedded Clustering (DiEC), a framework that performs unsupervised clustering by leveraging internal embeddings from the denoising network of a pretrained diffusion model. DiEC treats internal activations generated under different noise conditions and network hierarchies as a searchable representation trajectory, and builds the clustering pipeline on top of it: we select bottleneck activations as the Clustering Middle Layer (CML) and efficiently identify the clustering-optimal timestep (COT) via a fast scan on a data subset. We then introduce clustering optimization in this representation space, sharpening assignments with a $\mathrm{KL}$-divergence-based clustering loss, preserving local structure through adaptive graph embedding, and alleviating potential conflicts between denoising and clustering objectives via residual feature decoupling, which explicitly strengthens cluster structure and improves training stability. Experiments and systematic ablations show that DiEC outperforms existing deep clustering methods on multiple benchmark datasets.Our contributions are as follows:
  • We propose DiEC, the first framework to perform clustering by exploiting internal representations of pretrained diffusion models.
  • We introduce a representation-trajectory perspective and show that clusterability varies systematically over the timestep $\times$ layer space; based on this insight, we identify a more clustering-friendly layer location and develop an efficient procedure to quickly localize the clustering-optimal timestep.
  • We design a clustering-aware optimization that combines residual feature decoupling with a $\mathrm{KL}$-based clustering loss and graph-embedding regularization, improving training stability and mitigating conflicts between denoising and clustering objectives.
  • Extensive benchmark results demonstrate that DiEC achieves clear improvements on clustering metrics over deep clustering methods built on traditional generative representations such as AE/VAE/GAN.
2 Related Work
Clustering in Learned Embedding Spaces.
Most prior work follows a two-stage paradigm: it first learns a feature extractor $f(x)\rightarrow z$ and then performs cluster-assignment optimization in the embedding space $z$, where autoencoders and other generative models mainly serve as representation learners. An early representative method is DEC xie2016dec, which builds on a pretrained autoencoder and learns soft cluster assignments in a low-dimensional embedding space by matching them to an auxiliary target distribution. While DEC can learn clustering-oriented representations effectively, its performance is often sensitive to initialization, and training may suffer from degenerate solutions.
Building on variational autoencoders (VAEs), VaDE jiang2017vade explicitly injects a clustering structure into the latent space and jointly learns the generative model, latent representations, and cluster assignments under a variational inference framework. However, VaDE typically relies on a well-initialized pretrained stage, and can be sensitive to modeling assumptions and hyperparameters in high-dimensional or complex data regimes.
Within the GAN framework, ClusterGAN mukherjee2019clustergan introduces an additional inverse mapping $E(\cdot)$ alongside the generator, so that the inferred code $z=E(x)$ aligns with the latent prior of the generator. Training is guided by a closed-loop latent regression constraint, $z\rightarrow x=G(z)\rightarrow\hat{z}=E(x)$, encouraging $z$ to serve as a faithful generative-space representation from which cluster indicators can be read out (e.g., via discrete latent components) or obtained by clustering in $z$. However, such methods rely on adversarial min–max optimization and often face challenges in stability and scalability. In contrast, recent work has increasingly explored diffusion models as an alternative paradigm for generative modeling and representation learning ho2020ddpm; song2021scorebased.
Moreover, generic deep neural networks (DNNs), e.g., CNNs, are also widely used as feature extractors. These methods typically learn representations via contrastive learning or consistency regularization, and then optimize cluster assignments in the learned embedding space, with representative examples including Contrastive Clustering (CC) li2021cc and SCAN vangansbeke2020scan. Since this paper focuses on generative representation learning paradigms, we primarily compare clustering methods built on autoencoders and generative models; approaches based on contrastive learning or consistency constraints are not treated as our main comparison targets.
Diffusion Models as Feature Extractors.
In recent years, a growing body of work has explored using diffusion models as feature extractors to provide discriminative representations for downstream vision tasks; for a systematic overview, we refer to Fuest et al. fuest2024survey. Existing studies that leverage diffusion models for representation extraction can be broadly categorized into two groups.
The first line of work treats diffusion models as a representation source: without substantially modifying the model architecture, it directly reads intermediate features from the denoising network at selected timesteps or layers and uses them for downstream tasks. These methods typically use a pretrained diffusion model as a frozen or lightly adapted backbone, and focus on characterizing the discriminative power of features across timesteps and hierarchies. Representative examples include DDAE xiang2023ddae, DIFT tang2023dift, and Diffusion Hyperfeatures luo2023hyperfeatures.
A second line of work redesigns diffusion models as explicit representation learners by introducing structured latent variables, bottlenecks, or reconstruction constraints, so that diffusion training directly optimizes a semantically meaningful representation space instead of merely reading intermediate activations. Representative methods include Diffusion Autoencoders preechakul2022diffae, InfoDiffusion wang2023infodiffusion, and SODA hudson2024soda, which leverage the diffusion process itself to learn stable and decodable representations.
However, these approaches typically evaluate representation quality using discriminative criteria such as linear-probe classification, and do not explicitly address the challenges specific to fully unsupervised clustering.
Diffusion for Clustering.
Research on introducing diffusion models into fully unsupervised clustering remains limited. CLUDI uziel2025cludi trains a conditional diffusion model on top of frozen pretrained ViT features, samples assignment embeddings, and improves stability by averaging multiple samples; its core training signal is a teacher–student consistency alignment, while inference relies on multi-step reverse diffusion sampling and can be time-consuming. ClusterDDPM yan2025clusterddpm instead uses conditional diffusion to model cluster-related generative factors, making it closer to conditional generative modeling. In contrast, DiEC asks whether the denoising network itself can serve as a feature extractor for clustering, rather than relying on a frozen external representation or using diffusion mainly for assignment sampling and consistency modeling.
3 Background
Diffusion models.Diffusion models have recently emerged as a leading paradigm for high-fidelity generative modeling. Compared with traditional generative models such as VAEs and GANs, they often offer a better trade-off between sample fidelity and training stability, and have achieved strong results in image synthesis, text-to-image generation, image editing, super-resolution, and conditional generation dhariwal2021diffusion; nichol2022glide; saharia2021sr3; saharia2022imagen.
The core idea of diffusion models is to cast generation as an iterative denoising process: they first define a forward diffusion process that gradually perturbs data into approximately isotropic Gaussian noise, and then learn a corresponding reverse denoising process that progressively transforms noise back to the data distribution.
In DDPM ho2020ddpm, given a data sample $x_{0}\sim q_{\mathrm{data}}(x)$, where $q_{\mathrm{data}}(x)$ denotes the true data distribution, the forward diffusion process is defined as a Markov chain:
$q(x_{t}\mid x_{t-1})=\mathcal{N}\!\left(x_{t};\,\sqrt{1-\beta_{t}}\,x_{t-1},\,\beta_{t}\mathbf{I}\right).$ (1)
where $\{\beta_{t}\}_{t=1}^{T}$ denotes the noise schedule and $t\in\{1,\ldots,T\}$ indexes the diffusion timestep. Let $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$. Then,
$x_{t}=\sqrt{\bar{\alpha}_{t}}\,x_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon,\qquad\epsilon\sim\mathcal{N}(0,\mathbf{I}),$ (2)
indicating that the noise level of a sample is governed by the timestep $t$.
The reverse process is modeled by a learnable conditional distribution:
$p_{\theta}(x_{t-1}\mid x_{t})=\mathcal{N}\!\left(x_{t-1};\,\mu_{\theta}(x_{t},t),\,\sigma_{t}^{2}\mathbf{I}\right),$ (3)
where
$\mu_{\theta}(x_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\!\left(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\,\epsilon_{\theta}(x_{t},t)\right),$ (4)
$\sigma_{t}^{2}=\tilde{\beta}_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\,\beta_{t}.$ (5)
In practice, the neural network $\epsilon_{\theta}(x_{t},t)$ predicts the forward-process noise $\epsilon$ and is trained with a simple mean-squared error objective:
$\mathcal{L}(\theta)=\mathbb{E}_{t,x_{0},\epsilon}\!\left[\left\lVert\epsilon-\epsilon_{\theta}(x_{t},t)\right\rVert_{2}^{2}\right].$ (6)
This objective can be viewed as a simplified form of the evidence lower bound (ELBO). At sampling time, starting from $x_{T}\sim\mathcal{N}(0,\mathbf{I})$, we iteratively apply $p_{\theta}(x_{t-1}\mid x_{t})$ to progressively denoise and obtain generated samples.
It is worth noting that DDPMs are formulated with discrete timesteps, while subsequent work has also developed continuous-time variants and provided a unified view of diffusion processes from the perspective of stochastic differential equations (SDEs) song2021scorebased. The noise-prediction network $\epsilon_{\theta}(x_{t},t)$ is typically implemented with a U-Net architecture ronneberger2015unet, and recent studies have explored using Vision Transformers (ViTs) as the denoising backbone dosovitskiy2021vit; peebles2023dit. In this paper, we adopt the discrete-time DDPM formulation and a U-Net noise predictor to stay consistent with our subsequent analysis on timestep selection.
DEC-style Clustering Objective. Deep Embedded Clustering (DEC) performs unsupervised clustering by jointly learning sample embeddings and cluster centroids xie2016dec. Given an embedding $z_{i}$ and centroids $\{\mu_{k}\}_{k=1}^{K}$, DEC defines the soft assignment via a Student’s $t$-distribution kernel:
$q_{ik}=\frac{\left(1+\lVert z_{i}-\mu_{k}\rVert^{2}/\alpha\right)^{-\frac{\alpha+1}{2}}}{\sum_{k^{\prime}}\left(1+\lVert z_{i}-\mu_{k^{\prime}}\rVert^{2}/\alpha\right)^{-\frac{\alpha+1}{2}}}.$ (7)
To emphasize high-confidence assignments and mitigate cluster-size imbalance, DEC constructs a target distribution $P=[p_{ik}]$ from $Q=[q_{ik}]$:
$p_{ik}=\frac{q_{ik}^{2}/\sum_{i}q_{ik}}{\sum_{k^{\prime}}\left(q_{ik^{\prime}}^{2}/\sum_{i}q_{ik^{\prime}}\right)}.$ (8)
DEC then minimizes the KL divergence $\mathrm{KL}(P\|Q)=\sum_{i}\sum_{k}p_{ik}\log\frac{p_{ik}}{q_{ik}}$ for self-training, progressively sharpening the soft assignments and improving clustering consistency.
4 Diffusion Embedded clustering
We denote by $x_{t,i}$ the observation of the $i$-th sample at diffusion timestep $t$, and by $x_{0,i}$ the corresponding clean input. Given an unlabeled dataset $\mathcal{D}=\{x_{0,i}\}_{i=1}^{N}$ and the number of clusters $K$, DiEC aims to learn clustering representations from intermediate activations of a pretrained diffusion model and to predict a cluster assignment $\hat{y}_{i}\in\{1,\ldots,K\}$ for each sample. DiEC first locates a clustering-friendly diffusion representation at a specific layer–timestep pair $(l^{*},t^{*})$. Conditioning on this location, it then jointly optimizes the model with a clustering objective and a structural regularizer, yielding stable and separable clustering embeddings. Figure 1 provides an overview of the framework, and additional details are presented in the following subsections.
4.1 Locating Clustering-Friendly Diffusion Representations
We view the internal activations produced by a pretrained diffusion denoising network at different noise timesteps $t$ and network layers $l$ as a representation trajectory. Representations at different locations $(t,l)$ can exhibit markedly different clustering quality, which we measure using an unsupervised scoring function $S(\cdot)$. Ideally, one would jointly select the optimal representation in the two-dimensional $(t,l)$ space; however, a brute-force search requires $\mathcal{O}(|T||L|)$ evaluations (each involving feature extraction and clustering), which is prohibitively expensive. Empirically, we further observe that, for most datasets, the optimal layer remains stable over a wide range of timesteps, while the optimal timestep varies only mildly across nearby layers (see Sec. 5). Motivated by this weak-coupling behavior, we approximately decompose the 2D selection into two stages: we first fix the middle-layer readout location for clustering (CML), denoted by $l^{*}$, and then search for the clustering-optimal timestep (COT) on that layer, denoted by $t^{*}$.
Clustering Middle Layer (CML). In the denoising network of a pretrained diffusion model, features at different layers capture semantic abstractions at different granularity: shallow layers tend to encode local textures and noise details, whereas deeper layers are closer to the noise-prediction objective near the output. We define the bottleneck layer between the U-Net downsampling and upsampling paths as the Clustering Middle Layer (CML), denoted by $l^{*}$. For a sample $x_{0,i}$, given its noisy observation $x_{t,i}$ at timestep $t$, we read the activation at layer $l^{*}$ and flatten it to obtain
$e_{i}(t)=h_{\theta}^{l^{*}}(x_{t,i},t).$ (9)
In our experiments, the performance differences induced by layer selection are generally small (see Sec. 5). Therefore, DiEC does not treat layer search as a primary optimization target; instead, we fix CML to the U-Net bottleneck to balance computational efficiency and reproducibility. We accordingly focus on localizing the clustering-optimal timestep (COT) and the subsequent clustering-aware optimization.
Clustering-Optimal Timestep (COT). Under the condition of a fixed layer $l^{*}$,we define an unsupervised clustering-quality score $S(t)$ to measure clusterability at timestep $t$. We then define the clustering-optimal timestep as
$t^{*}=\arg\max_{t\in\mathcal{T}}S(t),$ (10)
where $\mathcal{T}$ is the candidate set of timesteps. We refer to $t^{*}$ as the Clustering-Optimal Timestep (COT).
To reduce the randomness introduced by diffusion noising, we optionally evaluate $S(t)$ by repeating the sampling at the same timestep $t$ for a few trials and averaging the scores. We describe in the next subsection how to efficiently localize $t^{*}$ while maintaining reliable evaluation, via the proposed Optimal Timestep Search (OTS).
Optimal Timestep Search (OTS). We seek to solve $t^{*}=\arg\max_{t\in\mathcal{T}}S(t)$, where $S(t)$ is an unsupervised score that measures the clusterability of $\{e_{i}(t)\}$. A naive strategy evaluates $S(t)$ for every $t\in\mathcal{T}$, which incurs $\mathcal{O}(|\mathcal{T}|)$ complexity, grows linearly with $|\mathcal{T}|$, and requires feature extraction and one clustering run per evaluation. To this end, we propose Optimal Timestep Search (OTS), which combines subset-based approximation and noise averaging to obtain a controlled-variance estimate of $S(t)$, and further leverages an approximate unimodality assumption to localize $t^{*}$ with only logarithmically many evaluations.
Subset-based approximation and noise-averaged scoring. Let the full unlabeled dataset be $\mathcal{D}=\{x_{0,i}\}_{i=1}^{N}$. OTS samples a subset $\mathcal{D}_{s}\subset\mathcal{D}$ with $|\mathcal{D}_{s}|=m$. For each $x_{0,i}\in\mathcal{D}_{s}$, at a given timestep $t$ we repeat $R$ forward noising trials by drawing $\xi_{i}^{(r)}\sim\mathcal{N}(0,\mathbf{I})$ and constructing
$x_{t,i}^{(r)}=\sqrt{\bar{\alpha}_{t}}\,x_{0,i}+\sqrt{1-\bar{\alpha}_{t}}\,\xi_{i}^{(r)},\qquad r=1,\ldots,R,$ (11)
together with the corresponding intermediate embeddings
$e_{i}^{(r)}(t)=h_{\theta}^{l^{*}}\!\left(x_{t,i}^{(r)},t\right).$ (12)
We then reduce stochasticity via noise averaging:
$\bar{e}_{i}(t)=\frac{1}{R}\sum_{r=1}^{R}e_{i}^{(r)}(t).$ (13)
Finally, we run a lightweight clustering (e.g., $k$-means) on $\{\bar{e}_{i}(t)\}_{x_{0,i}\in\mathcal{D}_{s}}$ and compute the Calinski–Harabasz index (CHI), yielding a subset- and noise-averaged estimate
$\widehat{\mathrm{CHI}}_{m,R}(t)=\mathrm{CHI}\!\left(\{\bar{e}_{i}(t)\}_{x_{0,i}\in\mathcal{D}_{s}}\right).$ (14)
Lemma 1 (Consistency of noise averaging). For a fixed sample $x_{i}$ and timestep $t$, if $\mathbb{E}\!\left[\left\lVert e_{i}^{(r)}(t)\right\rVert\right]<\infty$, then as $R\to\infty$,
$\bar{e}_{i}(t)=\frac{1}{R}\sum_{r=1}^{R}e_{i}^{(r)}(t)\xrightarrow{\mathrm{a.s.}}\mathbb{E}_{\xi}\!\left[e(x_{i},\xi,t)\right].$ (15)
Proof sketch. Since $\{e_{i}^{(r)}(t)\}_{r=1}^{R}$ are i.i.d. with respect to the noise $\xi$, the almost sure convergence follows from the strong law of large numbers.
Lemma 2 (Convergence of the subset statistic, conditional on stable partition). Let $z_{i}(t)=\mathbb{E}_{\xi}[e(x_{i},\xi,t)]$. If at timestep $t$ the optimal $k$-means partition remains stable under local perturbations (e.g., the optimum is unique with a positive margin), then as $m\to\infty$ and $R\to\infty$,
$\widehat{\mathrm{CHI}}_{m,R}(t)\xrightarrow{\mathrm{a.s.}}\mathrm{CHI}\!\left(\{z_{i}(t)\}_{i=1}^{N}\right).$ (16)
That is, the OTS score estimator concentrates as both the subset size and the noise averaging increase.
Proof sketch. By Lemma 1, $\bar{e}_{i}(t)$ converges to $z_{i}(t)$. Under the partition-stability condition, $\mathrm{CHI}$ is determined by sample statistics such as cluster means and within-/between-cluster scatter, which satisfy the law of large numbers under sample averaging; the desired convergence of $\widehat{\mathrm{CHI}}_{m,R}(t)$ then follows via continuity of the mapping from these statistics to $\mathrm{CHI}$.
4.2 Clustering-Aware Optimization
In this section, conditioned on $t^{*}$, we extract intermediate embeddings $e_{i}$ from the U-Net CML and obtain clustering representations $z_{i}$ via a lightweight residual mapping. We jointly optimize a DEC-style KL self-training objective and a graph regularizer on $z_{i}$ to strengthen cluster structure, while introducing a standard denoising loss at a random timestep $t_{\mathrm{rand}}$ to maintain diffusion consistency and stabilize the representations.
Residual-decoupled embedding. After determining $t^{*}$ and $l^{*}$, for each sample $x_{i}$ we construct its noisy input $x_{t^{*},i}$ at timestep $t^{*}$ and extract the bottleneck activation from the U-Net:
$e_{i}=h_{\theta}^{l^{*}}(x_{t^{*},i},t^{*}).$ (17)
We introduce a lightweight residual mapping $g_{\phi}(\cdot)$ and obtain the final clustering representation as
$z_{i}=e_{i}+g_{\phi}(e_{i}).$ (18)
Here, $g_{\phi}$ is a small MLP that captures clustering-relevant factors from $e_{i}$. Empirically, this mapping improves the discriminability between clusters while maintaining the stability of the underlying diffusion features.
Intuitively, clustering and diffusion denoising optimize different objectives and may impose conflicting requirements on representations. The residual design helps preserve the local structure of the original features, while allowing the clustering loss to adjust the representation toward being more clusterable, thereby mitigating interference from the denoising objective.
Clustering objective. For each sample $x_{0,i}$, we apply forward noising at the fixed optimal timestep $t^{*}$ to obtain a noisy observation
$x_{t^{*},i}=\sqrt{\bar{\alpha}_{t^{*}}}\,x_{0,i}+\sqrt{1-\bar{\alpha}_{t^{*}}}\,\epsilon,\qquad\epsilon\sim\mathcal{N}(0,\mathbf{I}).$ (19)
To smooth the stochasticity introduced by noise, we independently sample $M$ noise realizations at $t^{*}$ for the same sample, yielding $\{x_{t^{*},i}^{(m)}\}_{m=1}^{M}$, and extract the corresponding representations $\{z_{i}^{(m)}\}_{m=1}^{M}$ via the residual module. We then average them to obtain a stabilized feature
$z_{i}=\frac{1}{M}\sum_{m=1}^{M}z_{i}^{(m)}.$ (20)
We initialize cluster centroids $\{\mu_{k}\}_{k=1}^{K}$ using $k$-means on $\{z_{i}\}_{i=1}^{N}$ and then adopt a DEC-style self-training procedure: given the current centroids, we compute the soft assignments $Q=[q_{ik}]$ and construct the target distribution $P=[p_{ik}]$, and minimize
$\mathcal{L}_{\mathrm{KL}}=\mathrm{KL}(P\|Q)=\sum_{i}\sum_{k}p_{ik}\log\frac{p_{ik}}{q_{ik}},$ (21)
which progressively sharpens high-confidence assignments and strengthens cluster structure.
Adaptive graph regularization. Optimizing only $\mathcal{L}_{\mathrm{KL}}$ may be insufficient to preserve local geometric consistency in the embedding space. We therefore introduce an adaptive graph regularizer. Concretely, given the current representations, we construct a $k$-NN neighborhood $\mathrm{NB}_{i}$ for each sample and learn a row-normalized affinity matrix $\mathbf{S}=[s_{ij}]$. We impose a local-consistency constraint on the soft assignments $Q$:
$\mathcal{L}_{\mathrm{Gr}}$ $=\sum_{i=1}^{N}\sum_{j\in\mathrm{NB}_{i}}\sum_{k=1}^{K}s_{ij}\bigl(q_{ik}-q_{jk}\bigr)^{2},$ (22)
s.t. $\sum_{j\in\mathrm{NB}_{i}}s_{ij}=1,\qquad s_{ij}\in[0,1].$
and add an entropy regularizer to avoid degenerate weights:
$\mathcal{L}_{\mathrm{En}}=\sum_{i=1}^{N}\sum_{j\in\mathrm{NB}_{i}}s_{ij}\log s_{ij}.$ (23)
Joint training with diffusion consistency. The clustering branch imposes constraints only at the fixed optimal timestep $t^{*}$. If we directly update the shared denoising network parameters using this signal alone, the denoising capability at other timesteps may degrade, undermining the overall stability of diffusion representations. To prevent this, we introduce an additional consistency branch at random timesteps. At each iteration, we sample $t_{\mathrm{rand}}\sim\mathcal{U}(\{1,\ldots,T\})$, apply the standard forward noising to obtain $x_{t_{\mathrm{rand}},i}$, and train the network with the same noise-prediction MSE objective as in the diffusion background, thereby preserving the generative consistency of the pretrained diffusion model. Concretely, the diffusion-consistency loss is
$\mathcal{L}_{\mathrm{Re}}=\mathbb{E}_{t_{\mathrm{rand}},\epsilon}\!\left[\frac{1}{N}\sum_{i=1}^{N}\left\lVert\epsilon-\epsilon_{\theta}\!\left(x_{t_{\mathrm{rand}},i},t_{\mathrm{rand}}\right)\right\rVert_{2}^{2}\right].$ (24)
4.3 Loss Function
We optimize the diffusion U-Net (together with the residual module) end-to-end with a weighted combination of objectives. The overall training objective is
$\mathcal{L}=\alpha\,\mathcal{L}_{\mathrm{KL}}+\beta\,\mathcal{L}_{\mathrm{Gr}}+\gamma\,\mathcal{L}_{\mathrm{En}}+\mathcal{L}_{\mathrm{Re}},$ (25)
where $\mathcal{L}_{\mathrm{KL}}$ is the DEC-style clustering loss, $\mathcal{L}_{\mathrm{Gr}}$ and $\mathcal{L}_{\mathrm{En}}$ are the adaptive graph regularizer and the entropy regularizer (see Sec. 4.2), and $\mathcal{L}_{\mathrm{Re}}$ is the diffusion-consistency denoising reconstruction loss at random timesteps (see Sec. 4.2). The scalars $\alpha$, $\beta$, and $\gamma$ are weighting hyperparameters that balance the relative contributions of the loss terms.
We fix the weight of $\mathcal{L}_{\mathrm{Re}}$ to $1$ to keep the loss scale consistent with the pretraining objective and to avoid the clustering signal at the fixed $t^{*}$ dominating the shared-parameter updates. In practice, $\gamma$ mainly prevents degenerate graph weights and is not sensitive to its exact value, so a small constant suffices. The complete training procedure and the alternating updates are summarized in Algorithm 1.
5 Experiment
Datasets. We evaluate DiEC on four widely used benchmarks for unsupervised clustering: MNIST, USPS, Fashion-MNIST, and CIFAR-10. MNIST and Fashion-MNIST each contain 70,000 grayscale images of resolution $28\times 28$. USPS contains 9,298 grayscale handwritten-digit images with an original resolution of $16\times 16$. CIFAR-10 consists of 60,000 color images of resolution $32\times 32$. All datasets have $10$ ground-truth classes. In our clustering setup, we treat all samples as unlabeled during training and set $K$ to the true number of classes only for evaluation when computing clustering metrics. To match the diffusion model input, all images are normalized as a preprocessing step.
Dataset Samples Resolution Classes
MNIST 70,000 $28\times 28$ 10
USPS 9,298 $16\times 16$ 10
Fashion-MNIST 70,000 $28\times 28$ 10
CIFAR-10 60,000 $32\times 32$ 10
Table 1: Datasets used in our experiments.
Evaluation metrics. We evaluate clustering performance using three standard metrics: clustering accuracy (ACC), normalized mutual information (NMI), and adjusted Rand index (ARI). These metrics quantify the agreement between predicted cluster assignments and ground-truth labels from complementary perspectives, including optimal label matching, information-theoretic consistency, and partition similarity with chance correction. Higher values indicate better clustering quality.
Method MNIST USPS Fashion-MNIST CIFAR-10
DEC 84.3 61.9 51.8 30.1
IDEC 88.1 76.1
DCEC 89.0 79.0 60.0
VaDE 94.4 56.6 62.9 26.7
ClusterGAN 95.0 63.0 28.1
DAC 97.8 52.2
DECRA 96.4 95.1 96.3
GammM-VAE 95.2 70.8
ClusterDDPM 97.7 70.5 30.5
DiEC (Ours) 99.5 98.2 80.1 60.1
Table 2: Clustering accuracy (ACC, %) on four benchmarks. ACC is computed after optimal label matching (Hungarian algorithm). Full metrics (ACC/NMI/ARI) will be reported in v2.
Results. Table 2 summarizes clustering accuracy (ACC) on MNIST, USPS, Fashion-MNIST, and CIFAR-10. Overall, DiEC achieves the best performance on all four benchmarks, demonstrating the effectiveness of exploiting pretrained diffusion internal representations for unsupervised clustering. In particular, DiEC yields strong gains over classical generative clustering baselines on MNIST/USPS and remains robust on the more challenging Fashion-MNIST and CIFAR-10, highlighting the benefit of the proposed representation selection (CML+COT) and the joint optimization with diffusion consistency.
6 Conclusion
We presented DiEC (Diffusion Embedded Clustering), a diffusion-based deep clustering framework that performs unsupervised clustering by directly exploiting internal representations of a pretrained diffusion U-Net. DiEC views diffusion activations across layer $\times$ timestep as a representation trajectory and identifies a clustering-friendly location by fixing the bottleneck layer (CML) and efficiently localizing the clustering-optimal timestep (COT) via OTS. Conditioning on $(l^{*},t^{*})$, we jointly optimize a DEC-style KL self-training objective together with adaptive graph and entropy regularization, while introducing a random-timestep denoising-consistency branch to preserve diffusion stability. Experiments on four standard benchmarks demonstrate that DiEC achieves competitive clustering accuracy and validates the importance of selecting appropriate diffusion representations for clustering.